OpenStudy (anonymous):

Solve for x. 5(2x - 1) = 6 x = 1/10 x = 11/10 x = 1/2

OpenStudy (anonymous):

expand the parentheses

OpenStudy (anonymous):

 Means : 5(2x)+5(-1)=10x-5 So we have : 10x-5=6 and : 10x=6+5 and : 10x=11 so : x=11/10

OpenStudy (anonymous):

gracias :D

OpenStudy (anonymous):

What is the solution set of the following equation? -8x + x + 15 = -7x + 12 Ø {1/7}

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

-8x + x = -7x

OpenStudy (anonymous):

is their an app for this

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

their isnt i just checked :/

OpenStudy (anonymous):

Peggy had three times as many quarters as nickels. She had $1.60 in all. How many nickels and how many quarters did she have? Which of the following equations could be used to solve the problem? n + 3n = 1.60 n + 3n = 160 5n + 75n = 1.60 5n + 75n = 160

OpenStudy (anonymous):

probably the third one

OpenStudy (anonymous):

a nickel = 5 cent a quarter = 25 cent ?

OpenStudy (anonymous):

so three quarters are = 75 cent

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're welcome ;)